# Fourier-RWKV: A Multi-State Perception Network for Efficient Image Dehazing

### v1

@@ IV-A Experimental settings
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
7.62 dB 29.98 dB 33.95 dB 32.91 dB 36.01 dB 35.74 dB 37.19 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
11.54 dB 34.49 dB 30.01 dB 34.38 dB 40.42 dB 40.29 dB 40.96 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
16.49 dB 28.10 dB 34.53 dB 33.31 dB 38.96 dB 38.31 dB 40.32 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
13.36 dB 26.95 dB 28.42 dB 31.65 dB 34.99 dB 33.41 dB 38.21 dB PSNR
(a) Hazy image (b) FFA-Net [qin2020ffa] (c) MAXIM-2S [tu2022maxim] (d) Dehamer [9879191] (e) FocalNet [10377428] (f) OKNet [cui2024omni] (g) Ours (h) GT
Figure 5: Visual comparisons on synthetic hazy images from the SOTS dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
IV-B Comparison with State-of-the-art Methods
@@
In this section, we compare Fourier-RWKV with a wide range of state-of-the-art dehazing approaches, including CNN-based methods (GridDehazeNet [Liu_2019_ICCV], MSBDN [dong2020multi], FFA-Net [qin2020ffa], PMNet [ye2022perceiving], MAXIM-2S [tu2022maxim], DEA-Net [10411857], FocalNet [10377428], OKNet [cui2024omni], CSNet [cui2024hybrid], PGH2Net [su2025prior]), recent Transformer-based models (DeHamer [9879191], DehazeFormer [10076399], Fourmer [zhou2023fourmer], MB-TaylorFormer-B [10378631]), and a Mamba-based method (MAIR [DBLPconfmmFangZWZCH22]). Both quantitative and qualitative results are discussed below.
IV-B1 Quantitative Comparisons
@@
The quantitative comparison of Fourier-RWKV against SOTA models is summarized in Table I, where the best and second-best results are indicated in bold and underlined, respectively. Fourier-RWKV demonstrates outstanding performance across all test sets, with significant improvements in both PSNR and SSIM metrics.
@@
For the synthetic datasets, Fourier-RWKV achieves the second-best overall performance on the SOTS-Indoor subset. Its PSNR is 0.1 dB lower than PGH2Net, which is a similarly lightweight model. This minor gap primarily stems from PGH2Net’s integration of three distinct haze priors, giving it an advantage in relatively uniform synthetic indoor scenes. In comparison with another linear attention model MAIR [DBLPconfmmFangZWZCH22], Fourier-RWKV adds only 1.91 M additional parameters, while reducing FLOPs to 65.29% of MAIR, highlighting its exceptional computational efficiency. Despite a slight SSIM decrease of 0.001, Fourier-RWKV improves PSNR by 2.15 dB, underscoring its ability to recover fine image details and preserve structural integrity. On the SOTS-Outdoor subset, Fourier-RWKV outperforms all other models with a 2.05 dB PSNR gain and a 0.001 SSIM increase, reinforcing its robustness in handling more challenging outdoor haze scenarios.
@@
For the real-world datasets, Fourier-RWKV achieves the best performance across all evaluation metrics on both Dense-Haze [ancuti2019dense] and NH-HAZE [9150807]. Notably, on the NH-HAZE dataset, which focuses on non-uniform haze, our method surpasses the previous best results, improving PSNR by $0.35$ dB and SSIM by $0.03$, showcasing its strong generalization capability in complex real-world haze scenarios.
@@
IV-B2 Qualitative Comparisons
Fig. 5 shows the dehazing results on synthetic images from the SOTS dataset. Most comparative methods exhibit similar trends in both indoor and outdoor scenes. Earlier methods, such as FFA-Net [qin2020ffa], MAXIM-2S [tu2022maxim], and DeHamer [9879191], introduce edge blurring and artifacts in regions with significant local variations (Line 1, (a)–(c)) and struggle to fully remove haze in areas with large depth of field (Lines 2–3, (a)–(c)). In contrast, FocalNet [10377428] and OKNet [cui2024omni] show improved performance, while our method provides the most thorough dehazing and delivers the highest fidelity, especially in detail-rich regions.
Fig. 6 presents the dehazing results for two real-world scenes from the Dense-Haze [ancuti2019dense] dataset. In the first scene with extremely dense haze, the zoomed-in regions clearly show that Fourier-RWKV outperforms DeHamer [9879191] and MB-TaylorFormer-B [10378631] in detail restoration, while better preserving depth structure and spatial hierarchy. In the second scene, although our PSNR is slightly lower than that of DeHamer, our method excels in color fidelity and structural detail reconstruction, resulting in a higher SSIM score and better overall visual quality.
Fig. 7 compares two non-uniform real-world haze scenes from the NH-HAZE [9150807] dataset. The zoomed-in regions from both scenes reveal that Fourier-RWKV achieves the most comprehensive dehazing, effectively handling regions with large depth and complex textures, while recovering texture details more effectively than other methods.
@@ IV-C Ablation Studies
TABLE II: Ablation study for three core modules.
@@
Baseline 38.39 0.994 4.75 13.25
$\rightarrow$ DQ-Shift 39.86 0.995 4.84 14.03
@@
$\rightarrow$ DQ-Shift &
       Fourier Mix
40.82 0.996 5.29 15.58
@@
IV-C1 Effectiveness of Core Modules
To isolate the individual effects and interactions among modules, we design the following configurations. Starting with a baseline that includes Q-Shift, Spatial Mix, and $1\times 1$ convolutional skip connections, we construct six variants: (1) Baseline; (2) DQ-Shift replacing Q-Shift; (3) Fourier Mix replacing Spatial Mix; (4) SBM replacing $1\times 1$ convolutions; (5) Combined DQ-Shift and Fourier Mix replacing their respective counterparts; (6) Full model incorporating all three proposed modules. The results are presented in Table II.
As shown in Table II, the baseline achieves a PSNR of 38.39 dB on the SOTS-Indoor dataset. Replacing the corresponding components with DQ-Shift and Fourier Mix increases the PSNR by 1.47 dB and 1.90 dB, respectively, validating the effectiveness of spatial deformable perception and frequency-domain global perception. Combining these two modules raises the PSNR to 40.82 dB, underscoring the complementarity of the two perceptual states, which synergistically enhance the model’s degradation modeling capability.
@@ IV-C1 Effectiveness of Core Modules
The full model, integrating all three modules, achieves the best performance. Compared to the baseline, the parameter count and FLOPs increase by only 0.56 M and 2.44 G, respectively. This demonstrates that our proposed Multi-State Perception paradigm significantly boosts dehazing performance with relatively low overhead through the collaborative modeling of spatial, frequency-domain, and semantic relationships.
TABLE III: Ablation study for key component in DQ-Shift.
@@
TABLE IV: Ablation study for key component in Fourier Mix.
@@
TABLE V: Ablation study for key component in SBM.
@@
IV-C2 Effectiveness of DQ-Shift
To evaluate the impact of DQ-Shift, we design four variants: (1) The original fixed-pattern quad-directional shift; (2) A variant incorporating only the dynamic offset mechanism; (3) A combination of fixed and dynamic offsets, with the gating mechanism omitted; (4) Full DQ-Shift, which integrates both fixed and gated dynamic offsets. The results are reported in Table III.
As shown in Table III, compared to the full DQ-Shift, using only fixed or dynamic offset leads to a PSNR decrease of 0.83 dB and 0.32 dB, respectively. This highlights the importance of combining both offsets for optimal performance, with dynamic offset making a more significant contribution. When both offsets are used without the gating mechanism, PSNR drops by 0.24 dB, indicating that the gating mechanism effectively adjusts the sampling positions of dynamic offset, improving local modeling accuracy. Additionally, while this variant has slightly fewer parameters and FLOPs than the one using only dynamic offset, it achieves a PSNR improvement of 0.08 dB, showing that the fixed offset provides valuable directional information for spatial perception.
IV-C3 Effectiveness of Fourier Mix
To analyze the role of each component in Fourier Mix, we conduct ablation studies with four configurations: (1) Classic row scanning applied to spectral features; (2) Spectral sequencing/inversion with gating applied only in the spatial domain; (3) Spectral sequencing/inversion with gating applied only in the Fourier domain; (4) Full Fourier Mix, the complete implementation with dual-domain gating. The results are presented in Table IV.
As shown in Table IV, compared to the full Fourier Mix, using classical row-wise scanning of spectral features results in a PSNR drop of 0.49 dB, indicating that our proposed distance-based sorting strategy better preserves the relative relationships between frequency points, enabling more accurate global dependency modeling. Retaining only the spatial-domain gating causes a PSNR decrease of 1.08 dB, highlighting that the absence of Fourier-domain gating severely limits the global representation ability. Additionally, retaining only the Fourier-domain gating leads to a PSNR 0.46 dB lower than the full module, emphasizing the crucial role of spatial-domain gating in enhancing local spatial perception. These results collectively confirm the effectiveness and necessity of the proposed dual-domain gating mechanism.
IV-C4 Effectiveness of SBM
To validate the design rationale behind SBM, we develop five variants: (1) Replacing dynamic multi-scale semantic kernels with randomly initialized kernels of the same size; (2) Unifying multi-scale dynamic kernels into single-scale $5\times 5$ kernels; (3) Substituting KSFU with direct summation; (4) Adding semantic features to encoder features during feature correction, instead of replacing the DC component; (5) Full SBM, which integrates dynamic multi-scale semantic kernels, adaptive kernel selection fusion, and semantic replacement. The results are summarized in Table V.
As shown in Table V, compared to the full SBM, replacing dynamic multi-scale semantic kernels with randomly initialized kernels causes a 0.57 dB PSNR drop, demonstrating that kernels generated by the semantic similarity matrix more effectively capture semantic content. Using single-scale kernels results in a 0.14 dB PSNR decrease, emphasizing the advantage of multi-scale design in handling features of varying granularity. Replacing the KSFU module with direct summation leads to a 0.39 dB PSNR reduction, further confirming the benefits of adaptive kernel selection and fusion in cross-scale feature integration. Finally, adding semantic features instead of replacing the DC component causes a 0.27 dB PSNR decrease, suggesting that the semantic replacement strategy better introduces high-level semantic guidance while avoiding irrelevant low-frequency interference.
V Conclusion

### v2

@@ 4.1 Experimental settings
4.2 Comparison with State-of-the-art Methods
@@
In this section, we compare Fourier-RWKV with a wide range of state-of-the-art dehazing approaches, including CNN-based methods (GridDehazeNet [34], MSBDN [14], FFA-Net [43], PMNet [56], MAXIM-2S [47], DEA-Net [8], FocalNet [11], OKNet [12], CSNet [10], PGH2Net [46]), recent Transformer-based models (DeHamer [23], DehazeFormer [45], Fourmer [62], MB-TaylorFormer-B [44]), and a Mamba-based method (MAIR [30]). Both quantitative and qualitative results are discussed below.
Refer to caption Refer to caption Refer to caption Refer to caption
7.62 dB 29.98 dB 33.95 dB 32.91 dB
Refer to caption Refer to caption Refer to caption Refer to caption
11.54 dB 34.49 dB 30.01 dB 34.38 dB
(a) Hazy image (b) FFA-Net [43] (c) MAXIM-2S [47] (d) Dehamer [23]
Refer to caption Refer to caption Refer to caption Refer to caption
36.01 dB 35.74 dB 37.19 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption
40.42 dB 40.29 dB 40.96 dB PSNR
(e) FocalNet [11] (f) OKNet [12] (g) Ours (h) GT
Figure 5: Visual comparisons on synthetic hazy images from the SOTS-Indoor dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
4.2.1 Quantitative Comparisons
@@
The quantitative comparison of Fourier-RWKV against SOTA models is summarized in Table 1, where the best and second-best results are indicated in bold and underlined, respectively. Fourier-RWKV demonstrates outstanding performance across all test sets, with significant improvements in both PSNR and SSIM metrics.
Refer to caption Refer to caption Refer to caption Refer to caption
16.49 dB 28.10 dB 34.53 dB 33.31 dB
Refer to caption Refer to caption Refer to caption Refer to caption
13.36 dB 26.95 dB 28.42 dB 31.65 dB
(a) Hazy image (b) FFA-Net [43] (c) MAXIM-2S [47] (d) Dehamer [23]
Refer to caption Refer to caption Refer to caption Refer to caption
38.96 dB 38.31 dB 40.32 dB PSNR
Refer to caption Refer to caption Refer to caption Refer to caption
34.99 dB 33.41 dB 38.21 dB PSNR
(e) FocalNet [11] (f) OKNet [12] (g) Ours (h) GT
Figure 6: Visual comparisons on synthetic hazy images from the SOTS-Outdoor dataset. Key regions highlighted by red boxes are enlarged in the lower-left corner for clearer comparison.
@@
For the synthetic datasets, Fourier-RWKV achieves the second-best overall performance on the SOTS-Indoor subset. Its PSNR is 0.1 dB lower than PGH2Net, which is a similarly lightweight model. This minor gap primarily stems from PGH2Net’s integration of three distinct haze priors, giving it an advantage in relatively uniform synthetic indoor scenes. In comparison with another linear attention model MAIR [30], Fourier-RWKV adds only 1.91 M additional parameters, while reducing FLOPs to 65.29% of MAIR, highlighting its exceptional computational efficiency. Despite a slight SSIM decrease of 0.001, Fourier-RWKV improves PSNR by 2.15 dB, underscoring its ability to recover fine image details and preserve structural integrity. On the SOTS-Outdoor subset, Fourier-RWKV outperforms all other models with a 2.05 dB PSNR gain and a 0.001 SSIM increase, reinforcing its robustness in handling more challenging outdoor haze scenarios.
@@
For the real-world datasets, Fourier-RWKV achieves the best performance across all evaluation metrics on both Dense-Haze [1] and NH-HAZE [2]. Notably, on the NH-HAZE dataset, which focuses on non-uniform haze, our method surpasses the previous best results, improving PSNR by $0.35$ dB and SSIM by $0.03$, showcasing its strong generalization capability in complex real-world haze scenarios.
@@
4.2.2 Qualitative Comparisons
Figure 5 and Figure 6 show the dehazing results on synthetic images from the SOTS-Indoor and SOTS-Outdoor datasets, respectively. Most comparative methods exhibit similar trends in both indoor and outdoor scenes. Earlier methods, such as FFA-Net [43], MAXIM-2S [47], and DeHamer [23], introduce edge blurring and artifacts in regions with significant local variations (Figure 5 Line 1, (a)–(c)) and struggle to fully remove haze in areas with large depth of field (Figure 5, Line 2, and Figure 6, Line 1, (a)–(c)). In contrast, FocalNet [11] and OKNet [12] show improved performance, while our method provides the most thorough dehazing and delivers the highest fidelity, especially in detail-rich regions.
Figure 7 presents the dehazing results for two real-world scenes from the Dense-Haze [1] dataset. In the first scene with extremely dense haze, the zoomed-in regions clearly show that Fourier-RWKV outperforms DeHamer [23] and MB-TaylorFormer-B [44] in detail restoration, while better preserving depth structure and spatial hierarchy. In the second scene, although our PSNR is slightly lower than that of DeHamer, our method excels in color fidelity and structural detail reconstruction, resulting in a higher SSIM score and better overall visual quality.
Figure 8 compares two non-uniform real-world haze scenes from the NH-HAZE [2] dataset. The zoomed-in regions from both scenes reveal that Fourier-RWKV achieves the most comprehensive dehazing, effectively handling regions with large depth and complex textures, while recovering texture details more effectively than other methods.
@@ 4.3 Ablation Studies
Table 2: Ablation study for three core modules.
@@
Baseline 38.39 0.994 4.31 13.25
$\rightarrow$ DQ-Shift 39.86 0.995 4.40 14.03
@@
$\rightarrow$ DQ-Shift &        Fourier Mix 40.82 0.996 5.29 15.58
@@
4.3.1 Effectiveness of Core Modules
To isolate the individual effects and interactions among modules, we design the following configurations. Starting with a baseline that includes Q-Shift, Spatial Mix, and $1\times 1$ convolutional skip connections, we construct six variants: (1) Baseline; (2) DQ-Shift replacing Q-Shift; (3) Fourier Mix replacing Spatial Mix; (4) SBM replacing $1\times 1$ convolutions; (5) Combined DQ-Shift and Fourier Mix replacing their respective counterparts; (6) Full model incorporating all three proposed modules. The results are presented in Table 2.
As shown in Table 2, the baseline achieves a PSNR of 38.39 dB on the SOTS-Indoor dataset. Replacing the corresponding components with DQ-Shift and Fourier Mix increases the PSNR by 1.47 dB and 1.90 dB, respectively, validating the effectiveness of spatial deformable perception and frequency-domain global perception. Combining these two modules raises the PSNR to 40.82 dB, underscoring the complementarity of the two perceptual states, which synergistically enhance the model’s degradation modeling capability.
@@ 4.3.1 Effectiveness of Core Modules
The full model, integrating all three modules, achieves the best performance. Compared to the baseline, the parameter count and FLOPs increase by only 1.0 M and 2.44 G, respectively. This demonstrates that our proposed Multi-State Perception paradigm significantly boosts dehazing performance with relatively low overhead through the collaborative modeling of spatial, frequency-domain, and semantic relationships.
Table 3: Ablation study for key component in DQ-Shift.
@@
4.3.2 Effectiveness of DQ-Shift
To evaluate the impact of DQ-Shift, we design four variants: (1) The original fixed-pattern quad-directional shift; (2) A variant incorporating only the dynamic offset mechanism; (3) A combination of fixed and dynamic offsets, with the gating mechanism omitted; (4) Full DQ-Shift, which integrates both fixed and gated dynamic offsets. The results are reported in Table 3.
As shown in Table 3, compared to the full DQ-Shift, using only fixed or dynamic offset leads to a PSNR decrease of 0.83 dB and 0.32 dB, respectively. This highlights the importance of combining both offsets for optimal performance, with dynamic offset making a more significant contribution. When both offsets are used without the gating mechanism, PSNR drops by 0.24 dB, indicating that the gating mechanism effectively adjusts the sampling positions of dynamic offset, improving local modeling accuracy. Additionally, while this variant has slightly fewer parameters and FLOPs than the one using only dynamic offset, it achieves a PSNR improvement of 0.08 dB, showing that the fixed offset provides valuable directional information for spatial perception.
Table 4: Ablation study for key component in Fourier Mix.
@@ 4.3.2 Effectiveness of DQ-Shift
4.3.3 Effectiveness of Fourier Mix
To analyze the role of each component in Fourier Mix, we conduct ablation studies with four configurations: (1) Classic row scanning applied to spectral features; (2) Spectral sequencing/inversion with gating applied only in the spatial domain; (3) Spectral sequencing/inversion with gating applied only in the Fourier domain; (4) Full Fourier Mix, the complete implementation with dual-domain gating. The results are presented in Table 4.
As shown in Table 4, compared to the full Fourier Mix, using classical row-wise scanning of spectral features results in a PSNR drop of 0.49 dB, indicating that our proposed distance-based sorting strategy better preserves the relative relationships between frequency points, enabling more accurate global dependency modeling. Retaining only the spatial-domain gating causes a PSNR decrease of 1.08 dB, highlighting that the absence of Fourier-domain gating severely limits the global representation ability. Additionally, retaining only the Fourier-domain gating leads to a PSNR 0.46 dB lower than the full module, emphasizing the crucial role of spatial-domain gating in enhancing local spatial perception. These results collectively confirm the effectiveness and necessity of the proposed dual-domain gating mechanism.
Table 5: Ablation study for key component in SBM.
@@ 4.3.3 Effectiveness of Fourier Mix
4.3.4 Effectiveness of SBM
To validate the design rationale behind SBM, we develop five variants: (1) Replacing dynamic multi-scale semantic kernels with randomly initialized kernels of the same size; (2) Unifying multi-scale dynamic kernels into single-scale $5\times 5$ kernels; (3) Substituting KSFU with direct summation; (4) Adding semantic features to encoder features during feature correction, instead of replacing the DC component; (5) Full SBM, which integrates dynamic multi-scale semantic kernels, adaptive kernel selection fusion, and semantic replacement. The results are summarized in Table 5.
As shown in Table 5, compared to the full SBM, replacing dynamic multi-scale semantic kernels with randomly initialized kernels causes a 0.57 dB PSNR drop, demonstrating that kernels generated by the semantic similarity matrix more effectively capture semantic content. Using single-scale kernels results in a 0.14 dB PSNR decrease, emphasizing the advantage of multi-scale design in handling features of varying granularity. Replacing the KSFU module with direct summation leads to a 0.39 dB PSNR reduction, further confirming the benefits of adaptive kernel selection and fusion in cross-scale feature integration. Finally, adding semantic features instead of replacing the DC component causes a 0.27 dB PSNR decrease, suggesting that the semantic replacement strategy better introduces high-level semantic guidance while avoiding irrelevant low-frequency interference.
5 Conclusion